OpenStudy (anonymous):

6x / sqrt(-x-1) rationalize the denominator

OpenStudy (anonymous):

i got to -6x sqrt(-x-1) / -x-1 but i'm not sure how to simplify it after that

OpenStudy (tylerd):

this may involve imaginary numbers

OpenStudy (xapproachesinfinity):

multiply by \(\huge \frac{\sqrt{-x-1}}{\sqrt{-x-1}}\)

OpenStudy (xapproachesinfinity):

that should be root (-x-1) on top

OpenStudy (xapproachesinfinity):

darn it latex lol

OpenStudy (anonymous):

it would be \[-6\sqrt{x-1}\div-x-1\]

OpenStudy (anonymous):

but does it still need simplified further?

OpenStudy (xapproachesinfinity):

oh i didn't see you did that! not what else can you do xD

OpenStudy (tylerd):

has your teacher gone over imaginary numbers?

OpenStudy (anonymous):

\[6x \sqrt{x-1}\div-x-1\] but i know 6x and -x are like terms so i didn't know if I should divide them. and yes, but that's not what we're doing right now. it doesn't need simplified quite that far.

OpenStudy (xapproachesinfinity):

No no further operations need! that's the answer @TylerD, what's the need for imaginary here! I don't think so

OpenStudy (xapproachesinfinity):

needed*

OpenStudy (anonymous):

thanks!

OpenStudy (xapproachesinfinity):

yw!